solve the equation 3cos A + 2sin A = 2.8 for 0 degrees
Answer by stanbon(75887)

solve the equation 3cos A + 2sin A = 2.8
-------
3cos(A) + 2sqrt(1-cos^2(A)) = 2.8
-------------------------------------
2sqrt(1-cos^2(A)) = 2.8-3cos(A)
----
Square both sides to get:
4-4cos^2(A)) = 2.8^2 - 16.8cosA + 9cos^2(A)
----
13cos^2(A) - 16.8cos(A) + 3.84 = 0
------------------------------
cos(A) = [16.8 +- sqrt(16.8^2 -4*9*3.84)]/26
cos(A) = [16.8+-12]/26
cos(A) = 4.8/26 = 0.1846
A = arccos(0.1846) = 79.36 degrees
-----------------------------------------
Cheers,
Stan H.
==============